C-H STRETCH OVERTONE SPECTRA OF FLUORINATED ETHERS

\underline{SHIZUKA HSIEH}, Chemistry Department, Smith College, Northampton, MA 01063.

Photoacoustic spectra of some fluorinated ethers at 4-6 quanta of C-H stretch illustrate effects of fluorination on overtone frequencies and intensities.